# Interpretability and Generalization Bounds for Learning Spatial Physics

### v1

@@ 4 Experiments
We examine six different types of ML models. We validate our theoretical results by comparing analytical expressions with experimental results. We then demonstrate that generalization can be inferred from data with careful cross-validation. Finally, we demonstrate the interpretation of model weights as the Green’s function operator. App. B details software, hardware, and hyperparameters. App.  D contains additional results on grid size.
To construct the datasets, we use the spaces described in Section 3.1. There are 25 different function classes considered: one piecewise linear dataset (FEM) and $p\in[1,8]$ for polynomial, sine, and cosine spaces. A model is trained on one dataset, and then evaluated on the other 24 datasets with the same $N_{grid}$. For each choice $(type,p,N_{\text{grid}})$, we generate 10,000 examples to represent the infinite data limit. Each dataset is normalized to have unit norm in $u$ to (1) minimize the effects of numerical ill-conditioning and round-off errors in the $p\geq 5$ datasets with extreme magnitudes, and (2) to make the mean squared errors (MSE) comparable across datasets.
4.1 Comparison of Empirical Errors to Analytical Theory
Figure 2: Comparison of numerical experiments to analytically derived ML parameters from Theorems 1 and 2 trained on polynomial spaces with $N_{grid}=22$. (Left) The observed error for the linear matrix matches predictions for lower order $p$. (Middle and right) For the finite difference model, the analytical assumptions overestimate the error, but the trend matches with observations.
The theoretical results can be used to compute predicted optimal parameters, allowing for a priori estimations of errors as shown in Fig. 2. The equation from the proof of Theorem 1 can analytically solve expressions of $w$ as in App. A.2. The equation from Theorem 2 can be used to analytically predict ${\bm{W}}(p)$ by symbolically computing ${\bm{A}}$ and ${\bm{U}}$ using the procedure in A.5. The observed qualitative behavior matches that predicted by the theorems.
4.2 OOD Dataset Generalization
(a) Finite Different: $w[\mathtt{FD2}({\bm{u}},\Delta x^{2})]={\bm{f}}$
@@ 4.2 OOD Dataset Generalization
Figure 3: Heatmap of evaluation error for different training and testing combinations across four model types. Each row represents one model trained the dataset labeled on the y axis, and each column shows the MSE on corrosending the test set. The dashed lines separate the function class families into blocks. Within the each block, the cells to the left of a cell are function subspaces. The color map represents the MSE on an entire test dataset on a log scale: sharp transitions in color values correspond to ×$\times 10^{10}$ increase in MSE. Note the blue lower triangular blocks in 3(b) and 3(c): these are cases where the model generalizes satisfying $\mathcal{L}(w,\mathcal{F}_{test})\lesssim\mathcal{L}(w,\mathcal{F}_{train})$ when $\mathcal{F}_{\text{test}}\subseteq\mathcal{F}_{\text{train}}$.
We conduct the training procedure for six types of models on datasets constructed with $N_{\text{grid}}=22$. We train each type of model from scratch using each of the 25 different datasets as the training dataset. After training, each model is then evaluated on the 24 other datasets that unseen during its training. This results in a $25\times 25$ grid of MSEs for each model, shown in Fig. 3 and 4 as heatmaps. The model with the best training MSE from 5 runs is chosen to its corresponding row in the heatamp. The implementation details and hyperparameters for all experiments are provided in App. B. Line plots of the same data with error bars over multiple seeds are included in App. C.
@@
As predicted by Theorem 1, the test error is observed in Fig. 3(a) to increase as the polynomial order of the training data increases. The error is confounded by both the error from the finite difference and the error arising from the $|w-k|$ term. The error increases by both generalizing across higher order functions (the former term), moving to the right in the heatmap, and training on higher order functions (the latter term), moving down in the heatmap. Fig. 2 and Fig. 14 in App. D further explore this error in more settings. The error for the FEM dataset is orders of magnitude higher as it does not satisfy smoothness requirements for finite differences.
@@
We turn to Theorem 2 to empirically verify its predictions of ${\bm{W}}$ in Fig. 3(b). We initialize $W_{0}=0$, and do not use regularization. For each training run, the test error rises sharply when the test data are no longer in a subspace of the training data. There are three clear lower-triangular blocks in the heatmap representing the three families of nested subspaces, providing evidence for generalization across many types of function subspaces. Increasing polynomial orders decrease the MSE on the Cosine and Sine datasets (forming pyramid structures in the top middle and right.) The piecewise linear function space along the top row contains the $p=1$ space exactly, and consistently yields low MSE in all models.
@@
We apply nonlinear parameterizations to the linear problem by training two different deep learning models, a deep linear model and a shallow neural network. Deep linear models are a tractable stepping stone between linear models and MLP (Saxe et al., 2013; Arora et al., 2018) to understand nonlinear training dynamics. Both models are trained with a larger hidden dimension than the input and output size (App. B). As seen in Figs. 3(c) and 3(d), both models do not exhibit generalization behavior consistently. Both achieve low test error on $p_{\text{train}}$, but the test error sharply increases for any other dataset where $p_{\text{test}}\neq p_{\text{train}}$. The MLP is strongly diagonal, showing that the model will overfit to the training data, and not generalize in any case. The Deep Linear model shows mixed behavior: it exhibits subspace generalization on the sine and cosine datasets but not on the polynomial datasets. As with the linear model, some cross-subspace generalization is observed when the models are trained with piecewise linear functions.
@@
(a) DeepONet: $u(x)=trunk({\bm{f}})\cdot branch(x)$
(b) Neural Op.: $u(x)=\mathtt{fft}^{-1}(R_{\phi}\cdot\mathtt{fft}({\bm{f}}))$
We now investigate two models that are designed for learning PDEs: the DeepONet Lu et al. (2021a) and the Fourier Neural Operator Kovachki et al. (2021). The results are displayed in Fig. 4. The DeepONet and the Neural Operator support flexible $x$-coordinate distributions, but for both of these methods we still consider the regular grid, and do not consider grid adaptation of these methods. The DeepONet shows block lower-triangular structure similar to that of the linear model. The diagonal bands (evaluation on the training distribution) exhibit lower error values, showing a slight overfitting that does not strictly satisfy the subspace generalization. The dips in the trend are clear in Fig. 11 in App. C, where the model could satisfy $\mathcal{L}_{test}\lesssim\mathcal{L}_{train}$ tolerance depending on the application. The Neural Operator has a similar generalization pattern, but did not achieve a low training MSE on all higher order function classes.
4.3 Extracting Green’s functions from black-box models.
@@
Figure 6: Visualization of test function evaluations, $\text{Model}({\bm{e}}_{i})$, on black-box models can reveal Green’s function structure. The top row shows one test function evaluation at the node $x=\Delta x$ (each column has a different y-scale), and the bottom row shows a matrix constructed from model evaluations for all test functions. The colormap limits on the bottom heatmaps are set to match the y-scale of the corresponding plot above them.
@@
For the linear model, this is the same as inspecting the weight matrix, ${\bm{A}}_{ij}\leftrightarrow{\bm{W}}_{ij}$. Fig. 5 elaborates on Fig. 1 to show the convergence of the weight matrix to the derived analytical solution for increasing modes (or complexity) of data. For lower orders $p$, the extracted functions are inscrutable, even though the model achieves low error on the training set. Training the linear model on the FEM dataset is sufficient to recover an accurate Green’s operator ${\bm{A}}$. It is even possible to invert the weight matrix (${\bm{W}}^{-1}=\hat{{\bm{L}}}$) to uncover a local stencil structure, which is evident in its qualitatively tridiagonal appearance.
Fig. 6 shows that it is also possible to extract the Green’s functions for deep models by evaluating the nonlinear models on one-hot inputs ${\bm{f}}=\textbf{e}_{j}$. For $p=3$, there is no obvious evidence of the expected structure, even the model achieves low MSE on the training set. The expected structure is extracted using the MLP and DeepONet when they are trained on FEM data. The structure is somewhat evident for the Fourier Neural Operator with the $cos[7]$ training run, but no others.
@@ 5 Conclusion
Learning spatially dependent problems is limited by the underlying function space of the training data and the observation discretization. We theoretically proved that this barrier exists on the simplest problem — the Poisson problem — with the simplest models, parameter fitting finite differences and linear models. We illustrated that carefully choosing the training data can mitigate this issue and that it is possible to generalize to other function spaces, which can inform data collection and cross-validation techniques. As another form of validation, black-box models can be directly interpreted as integrated Green’s functions when they generalize effectively across many test domains. However, this is still pessimistic and presents a headwind for discovering unknown physics from real world data, given the inherent limitations in real world data collection.
We demonstrated this behavior across a wide range of representative model types, from simple parameter fits to models designed for PDE learning, but it may be possible to design a modeling technique or data augmentation strategy to mitigate this issue. We did not consider any methods that incorporate a PDE loss or physics-informed loss, such as Physics Informed Neural Networks (PINNs) or Physics Informed Neural Operators (PINOs), because we were specifically studying the case of discovering the PDE operator from data without any prior knowledge. Our analysis of the finite different fitting suggest that even the strongest physical priors may not mitigate the impact of the training data distribution. Even more confoundingly, different types ML models can exhibit different and sometimes opposing generalization behaviors.
While our theory required a priori knowledge of how the dataset was constructed, our results suggest that extracting generalized scientific knowledge may require new methods capable of compensating for these limitations by discovering the underlying subspace of the data. We propose this methodology to be an evaluation benchmark on future development of learning physics. As we have seen that even the Poisson equation can pose a challenge, we recommend that future proposals for PDE learning techniques measure cross-set generalization on Poisson data.
Limitations
This work is an initial theoretical exposition of the newly discovered behavior, and leaves many questions open. We only performed the exploration and proved theorems on a single differential equation to focus on demonstrating the result across different models. Future work should extend this analysis to more complex physical systems and ML models, and construct a similar evaluation on real-world data. We do not currently have any recommendations for how to solve the lack of generalization; we leave this challenge open to future research to utilize the proposed evaluation method to improve existing techniques and innovate for learning of physical systems.
@@ A.1 Proof of Theorem 1
$f^{(n)}(x)=-k\sum_{m=0}^{p}c_{m}x^{m},\quad u^{(n)}(x)=\sum_{m=-2}^{p}\frac{c_{m}x^{m+2}}{(m+2)(m+1)}$ (13)
for coefficients $c_{m}$. Recall that a finite difference stencil (LeVeque, 2007) of accuracy order $q$ will satisfy
@@
$FD_{q}(u,\Delta x)=\sum_{m=0}^{q-1}c_{m}x^{m}+\sum_{m=q}^{p}c_{m}(1+\zeta_{m}\Delta x^{m})x^{m}:=\sum_{m}b_{m}x^{m}$ (15)
where $\zeta_{m}$ is a bound of the coefficient of error introduced for each term $m$ starting at $q$ by Taylor’s theorem. These error terms are are dependent upon the finite difference stencil; the following section illustrates two examples. This proof is for the general case. For the usual 3 point stencil, $q=2$. Simplifying the loss from (29)
[…]
@@
[…]
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}(wb_{m}-kc_{m})(wb_{j}-kc_{j})\frac{1}{m+j+1}\right]$ (19)
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)c_{m}c_{j}\frac{1}{m+j+1}\right]$ (20)
$=\sum_{j=0}^{p}\sum_{m=0}^{p}(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)\frac{\mathbb{E}_{c}\left[c_{m}c_{j}\right]}{m+j+1}$ (21)
@@
Note that $w=k$ if $p<q$ as the numerator vanishes. Conversely, when $p>q$, $w\to k$ as $\Delta x\rightarrow 0$ at a rate of $\Delta x^{q}$ by polynomial division. Furthermore, for a fixed $\Delta x$, increasing $p$ Each additional term adding $p$ adds additional terms: at the very least, the term $\xi_{m}\Delta x^{2m}$ is always positive, so the magnitude of that term will always increase, even if $\zeta_{m}\Delta x^{m}$ has favorable canceling out.

### v2

@@ 4 Experiments
(a) FD: $w[\mathtt{FD2}({\bm{u}},\Delta x^{2})]={\bm{f}}$
@@
Figure 4: Cross-evaluation error heatmaps for four model types. Rows: training set; columns: test set. Dashed lines separate function class families; cells left of each cell are subspaces. Color shows log-scale MSE (sharp transitions $\approx\times 10^{10}$). The blue lower triangular blocks in 4(b) and 4(c) indicate generalization with $\mathcal{L}(w,\mathcal{F}_{test})\lesssim\mathcal{L}(w,\mathcal{F}_{train})$ when $\mathcal{F}_{\text{test}}\subseteq\mathcal{F}_{\text{train}}$.
We examine eight different types of ML models in different settings: parameter fitting inverse problems where the DE is known (finite difference, PINN), black-box models (linear, deep linear, multilayer perceptron), SciML black-box models (DeepONet, Neural Operator), and physics informed learning models (Physics Informed Deep ONet). The theoretical closed form error estimates are replicated empirically. We then demonstrate that generalization can be inferred from data with careful cross-validation. The results are replicated in the presence of noise and on different linear PDEs. Finally, we demonstrate that the Green’s function operator can be extracted from black-box models. App. B details software, hardware, and hyperparameters. App.  D contains additional results on grid size variation.
A total of 25 different datasets are constructed using the function spaces described in Section 3.1: one piecewise linear dataset (FEM) and $p\in[1,8]$ for polynomial, sine, and cosine spaces. For each choice $(type,p,N_{\text{grid}})$, we generate 10,000 examples to represent the infinite data limit. Each dataset is normalized to have unit norm in $u$ to (1) minimize the effects of numerical ill-conditioning and round-off errors in the $p\geq 5$ datasets with extreme magnitudes, and (2) make the MSE comparable across datasets.
4.1 Comparison of Empirical Errors to Analytical Theory
The theoretical results can be used to compute predicted optimal parameters, allowing for a priori estimations of errors as shown in Fig. 2. The result in Theorem 3.1 can analytically solve expressions of $w$ as in App. A.2. The result Theorem 3.2 can analytically predict ${\bm{W}}(p)$ using the procedure in App. A.4. The observed qualitative behavior matches that predicted by the theorems.
4.2 PINN Inverse Problem for Parameter Learning
We analyze inverse problem solving using PINNs using the DeepXDE library to learn parameter of the 1D Poisson problem, similar to the finite difference model. Vanilla PINNs are not equation discovery methods but are used as numerical solvers when both the PDE is known and some data points are known. The PDE loss can be naturally coupled unknown parameters in the PINN optimization to solve an inverse problem:
$\min_{\theta,w}\sum_{x\in grid}\left(w\frac{d^{2}}{dx^{2}}N(\theta,x_{i})+f(x_{i})\right)^{2}+\left(N(\theta_{,}x_{i})-u_{i}\right)^{2}$ (12)
where the neural network parameters $\theta$ and the unknown physical constant $w$ (which is expected to converge to the physical constant $k$) are learned simultaneously. The neural network $N(\theta,x_{i})\approx u(x)$ fits to a single solution instance $\{x_{i},f_{i},u_{i}\}$. The results in Fig. 3 match that of the finite difference white box method, where the parameter error wkw-k increases with increasing function space terms.
[…]
Figure 6: The generalization pattern for the Physics-Informed DeepONet is similar to the Deep ONet in Fig. 5, but with a higher baseline error on the training data.
4.3 OOD Dataset Generalization
We train models from scratch on datasets with $N_{\text{grid}}=22$ using each of the 25 different datasets. After training, each model is evaluated on the 24 other datasets unseen during training. This results in a $25\times 25$ grid of MSEs for each model, shown in Figs. 48 as heatmaps. The model with the best training MSE from 5 runs is chosen to plot the corresponding row in the heatamp. Error bars over multiple seeds are shown in App. C.
@@ 4.3 OOD Dataset Generalization
As predicted by Theorem 3.1, the test error is observed in Fig. 4(a) increases by both generalizing across higher order functions, moving to the right in the heatmap, and training on higher order functions, moving down in the heatmap. The FEM dataset error is orders of magnitude higher as it violates smoothness requirements for finite differences. See 13 in App. D for further analysis on $\Delta x$ terms.
@@
We turn to Theorem 3.2 to empirically verify its predictions of $\bm{W}$} in Fig. 4(b). We initialize $W_{0}=0$ and do not use regularization. For each training run, the test error rises sharply when the test data are no longer in a subspace of the training data. There are three clear lower-triangular blocks in the heatmap representing the three families of nested subspaces, providing evidence for generalization across function subspaces. Increasing polynomial orders decrease the MSE on the cosine and sine datasets (forming pyramid structures in the top middle and right.) The FEM piecewise linear function space along the top row show consistently low MSE in all models.
@@
We train two basic nonlinear models: a deep linear model and a shallow neural network. As seen in Figs. 4(c) and 4(d), neither models exhibits generalization behavior consistently. The MLP is strongly diagonal, showing that the model will overfit to the training data, and not generalize in any case. The Deep Linear model shows mixed behavior: it exhibits subspace generalization on the sine and cosine datasets but not on the polynomial datasets. As with the linear model, some cross-subspace generalization is observed when the models are trained with piecewise linear functions.
@@
We investigate two models that are designed for learning PDEs: the DeepONet Lu et al. (2021a) and the Fourier Neural Operator Kovachki et al. (2021). The results are displayed in Fig. 5. Both models support flexible $x$-coordinate distributions; testing in our experiments is still performed on the regular grid for consistency. The DeepONet shows block lower-triangular structure similar to that of the linear model. The diagonals (evaluation on the training distribution) exhibit lower error values, showing a slight overfitting that does not strictly satisfy the subspace generalization. The dips in the trend are clear in Fig. 11(e) in App. C, where the model could satisfy $\mathcal{L}_{test}\lesssim\mathcal{L}_{train}$ tolerance depending on the application. The Neural Operator has a similar generalization pattern, but did not achieve a low training MSE on all function classes.
Physics-Informed DeepONet
We implement a PI-DeepONet following Wang et al. (2021) (Fig. 6), which uses the loss:
$\min_{\theta}\sum_{f\in data}\sum_{x\in grid}\left(-k\frac{d^{2}}{dx^{2}}N(\theta,\mathbf{f},x_{i})-f(x_{i})\right)^{2}+\sum_{x\in boundary}(\text{BC Loss}(N(\theta,\mathbf{f},x_{i})))^{2}$ (13)
where $N(\cdot)$ is the factorized DeepONet that takes the observation point $x_{i}$ and the forcing function values $\mathbf{f}$ as inputs. Boundary conditions are enforced through an additional loss term. The baseline error is larger than the other models, achieving only $10^{-6}$ error. The subspace generalization pattern is still evident, although less pronounced due to the raised floor of the error, showing that physics informed losses do not mitigate the subspace limitation.
Effect of Noise
Figure 7: Poisson equation with training data corrupted by noise.
We investigate the effect of measurement noise on the generalization boundaries by adding normally distributed noise to the 1D Poisson equation training data of $u$ and $f$. The results are displayed in Fig. 7. Rather than “smearing” the sharp generalization boundaries, noise raises the floor of achievable losses, achieving a lowest error of $\approx 10^{-9}$ compared to $10^{-20}$ error, and the generalization pattern is still present.
2D Poisson and Biharmonic Equations
(a) 2D Poisson Equation
(b) Biharmonic Equation
Figure 8: Generalization to out-of-distribution test datasets constructed for the 2D Poisson and Biharmonic equations learned with a linear model.
We repeat the same experiments for the 1D biharmonic and 2D Poisson equations using a linear model, shown in Fig. 8. The biharmonic equation, $k\frac{d^{4}u}{dx^{4}}=f$, displays the same block lower-triangular structure as the 1D case. Solutions for the biharmonic are constructing the same way as the 1D Poisson equation. For the 2D Poisson equation, forcing functions are constructed using independent basis functions in $x$ and $y$, $f(x,y)=\sum_{i=1}^{p}c_{i}\sin(\pi ix)\sum_{j=1}^{q}d_{j}\sin(\pi jy)$. This exhibits subspace generalization independently in each spatial direction, producing the Sierpiński-triangle-esque pattern. This shows that the results hold generally for linear PDEs, as implied by Theorem 3.2.
4.4 Extracting Green’s functions from black-box models
@@
Figure 10: Visualization of test function evaluations, $\text{Model}({\bm{e}}_{i})$, on black-box models can reveal Green’s function structure (bottom row). The top row shows one test function evaluation at the node $x=\Delta x$, and the bottom row shows a matrix constructed from model evaluations for all test functions. The colormap scale matches the y-scale of the corresponding plot above them.
@@
For the linear model, this is the same as inspecting the weight matrix, ${\bm{A}}_{ij}\leftrightarrow{\bm{W}}_{ij}$. Fig. 9 elaborates on Fig. 1 to show the convergence of the weight matrix to the derived analytical solution for increasing modes (or complexity) of data. For lower orders $p$, the extracted functions are inscrutable, even though the model achieves low error on the training set. Training the linear model on the FEM dataset is sufficient to recover an accurate Green’s operator $\bm{A}$}. It is even possible to invert the weight matrix ($^{\bm{W}}^{-1}=\hat{{\bm{L}}}$) to uncover a local stencil structure, which is evident in its tridiagonal appearance.
Fig. 10 shows that it is also possible to extract the Green’s functions for deep models by evaluating the nonlinear models on one-hot inputs ${\bm{f}}=\textbf{e}_{j}$. For $p=3$, there is no apparent structure, even the model achieves low MSE on the training set. The expected structure is extracted using the MLP and DeepONet when they are trained on FEM data. The structure is somewhat evident for the Fourier Neural Operator with the $cos[7]$ training run, but no others.
@@ 5 Conclusion
Learning spatially dependent problems is limited by the underlying function space of the training data and the observation discretization. We theoretically proved that this barrier exists on the simplest problem — the Poisson problem — on parameter fitting finite differences and linear models. We illustrated that carefully choosing the training data can mitigate this issue and that it is possible to generalize to other function spaces, which can inform data collection and cross-validation techniques. As another form of validation, black-box models can be directly interpreted as integrated Green’s functions when they generalize effectively across many test domains. However, this is still pessimistic and presents a headwind for discovering unknown physics from real world data, given the inherent limitations in real world data collection. Confoundingly, different types of ML models can exhibit different and sometimes opposing generalization behaviors.
We demonstrated this behavior across a wide range of representative model types, from simple parameter fits to models designed for PDE learning. Physics informed losses which incorporate prior knowledge (e.g., PINNs and Physics Informed DeepONets) suffer the same subspace limitation as pure black-box models. Our analysis suggests that even the strongest physical priors may not mitigate the impact of the training data distribution.
While our theory required a priori knowledge of how the dataset was constructed, our results suggest that extracting generalized scientific knowledge may require new methods capable of compensating for these limitations by discovering the underlying subspace of the data. We propose this methodology to be an evaluation benchmark on future development of learning physics. As we have seen that even linear DEs can pose a challenge, we recommend that future proposals for DE learning techniques measure cross-set generalization.
Impact Statement
This paper analyzes limitations in discovering underlying physics from data across an entire spectrum of modeling approaches, with implications for evaluation and validation in scientific ML. These findings are relevant for applications in physics and engineering, where understanding out-of-distribution behavior is important for safe real-world deployment. The work does not introduce new capabilities with direct societal risk.
Acknowledgments
Placeholder for blind review.
@@ Appendix A Proofs
$f^{(n)}(x)$ $=-k\sum_{m=0}^{p}c_{m}x^{m}$ (15)
$u^{(n)}(x)$ $=\sum_{m=-2}^{p}\frac{c_{m}x^{m+2}}{(m+2)(m+1)}$ (16)
for coefficients $c_{m}$. Recall that a finite difference stencil of accuracy order $q$ will satisfy
@@
$FD_{q}(u,\Delta x)$ $=\sum_{m=0}^{q-1}c_{m}x^{m}+\sum_{m=q}^{p}c_{m}(1+\zeta_{m}\Delta x^{m})x^{m}$ (18)
$:=\sum_{m}b_{m}x^{m}$ (19)
where $\zeta_{m}$ is a bound of the coefficient of error introduced for each term $m$ starting at $q$ by Taylor’s theorem. These error terms are are dependent upon the finite difference stencil; the following section illustrates two examples. This proof is for the general case. For the usual 3 point stencil, $q=2$. Simplifying the loss
$\mathcal{L}$ $=\mathbb{E}_{c}\left[\int_{0}^{1}\left(w\sum_{m=0}^{p}b_{m}x^{m}-k\sum_{m=0}^{p}c_{m}x^{m}\right)^{2}dx\right]$ (20)
$=\mathbb{E}_{c}\left[\int_{0}^{1}\left(\sum_{m=0}^{p}(wb_{m}-kc_{m})x^{m}\right)^{2}dx\right]$ (21)
@@
$\mathcal{L}$ $=\mathbb{E}_{c}\left[\int_{0}^{1}\left(\sum_{j=0}^{p}\sum_{m=0}^{p}(wb_{m}-kc_{m})(wb_{j}-kc_{j})x^{m+j}\right)dx\right]$ (22)
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}\frac{(wb_{m}-kc_{m})(wb_{j}-kc_{j})}{m+j+1}\right]$ (23)
$=\mathbb{E}_{c}\left[\sum_{j=0}^{p}\sum_{m=0}^{p}\frac{(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)c_{m}c_{j}}{m+j+1}\right]$ (24)
$=\sum_{j=0}^{p}\sum_{m=0}^{p}\frac{(w(1+\xi_{m})-k)(w(1+\xi_{j})-k)\mathbb{E}_{c}\left[c_{m}c_{j}\right]}{m+j+1}$ (25)
@@
Note that $w=k$ if $p<q$ as the numerator vanishes. Conversely, when $p>q$, $w\to k$ as $\Delta x\rightarrow 0$ at a rate of $\Delta x^{q}$ by polynomial division. Furthermore, for a fixed $\Delta x$, increasing $p$ . Each additional term adding $p$ adds additional terms: at the very least, the term $\xi_{m}\Delta x^{2m}$ is always positive, so the magnitude of that term will always increase, even if $\zeta_{m}\Delta x^{m}$ has favorable canceling out.